What is the missing number; 0,1,1,2,3,5,8,13,_ 34, 55?
Answer by checkley77(12844)

0,1,1,2,3,5,8,13,(21),34, 55?
0+1=1
1+1=2
1+2=3
2+3=5
3+5=8
8+13=21
13+21=34
21+34=55